What mass of phosphorus is present in 25 g P4S3? A. 14g B. 4.5g C. 11g D. 3.2g E. 100g
Answer #1

Molar mass of P is 31

Molar mass of P4S3=a(4*mass of P)+(3*atomic mass of S)

= (4*31)+(3*32)

= 220 g/mole

220 g of P4S3 contains 4*31= 124 g of P

25 g of P4S3 contains (25g*124g)/220g= 14 g of P

Therefore option A is correct.